OpenStudy (anonymous):

Below are the steps to solve an equation: Step 1: |x - 4| + 1 = 8 Step 2: |x - 4| = 8 - 1 Step 3: |x - 4| = 7 Which of the following is a correct next step to solve the equation?

OpenStudy (anonymous):

A correct next step is to add 4 to both sides to get x = 11.

OpenStudy (anonymous):

I disagree.

OpenStudy (anonymous):

@RyleeC explain your reasoning then

OpenStudy (anonymous):

So okay

OpenStudy (anonymous):

Absolute value must always be solved inside then pulled out as positive. Therefore x-4 must be positive, but must be subtracted first. So you would have to minus 7 by 4, to get x = 3

OpenStudy (anonymous):

So would I do 7 -4

OpenStudy (anonymous):

I know, but let's say that x = 11, like I said earlier. If you take |11 - 4| = 7, the absolute value of 11 - 4 is 7.

OpenStudy (anonymous):

You can even go all the way back and do that. |11 - 4| + 1 = 8, 7 + 1 = 8, 8 = 8.

OpenStudy (anonymous):

But if x is -3 the absolute value would still be 7 since -3 - 4 is -7 but become +7 due to absolute value.

OpenStudy (anonymous):

SO would the next step be x + 4 = 7

OpenStudy (anonymous):

Ok, then. Let me solve it here (for is x = -3). |-3 + 4| + 1 = 8, |4 - 3| + 1 = 8, 1 + 1 = 8, 2 ≠ 8

OpenStudy (anonymous):

It works both ways x=-3, and 11. Therefore the correct step is to minus the 4 from both sides, to theoretically minus out the absolute value of a -4.

OpenStudy (anonymous):

You messed up your signs above^

OpenStudy (anonymous):

No, I just explained that -3 won't be an acceptable answer ^^^ The correct answer is x = 11, because you ADD 4, because on the left side, it's being SUBTRACTED.

OpenStudy (anonymous):

wht do u meen by meh messing up meh signs?

OpenStudy (anonymous):

You incorrectly explained the -3 because your signs are wrong.

OpenStudy (anonymous):

-3-4 is -7 which turns into 7 due to the absolute value and 7 = 7.

OpenStudy (anonymous):

The answeres are x - 4 = -7 x + 4 = -7 -x - 4 = 7 x + 4 = 7

OpenStudy (anonymous):

Its not -3 + 4.

OpenStudy (anonymous):

Oh, derp. So, x = -3 and x = 11.

OpenStudy (anonymous):

YES! ahaha

OpenStudy (anonymous):

There's multiple answers.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

Its D.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Welcome. :)

OpenStudy (anonymous):

Heres why. The technical correct way it to minus the 4, this takes the absolute value into consideration.

OpenStudy (anonymous):

D is the only answer that considers that step.